OpenStudy (anonymous):

Prove equation is an identity: Cos(x)Csc(x)Tan(x)=1

OpenStudy (anonymous):

\[\frac{ \cos (x) }{ 1 } \times \frac{ 1 }{ \sin(x) } \times \frac{ \sin (x) }{ \cos (x) }\] = 1

OpenStudy (anonymous):

I just need help Simplifying

OpenStudy (tkhunny):

See if it works for x = 0 Rewrite - multiplication is commutative. Swap the sine and cosine in the denominator. That will get you started.

OpenStudy (anonymous):

We're not supposed to solve for x?

OpenStudy (anonymous):

If I multiply \[\frac{ \cos (x) }{ \sin (x) } \times \frac{ \sin (x) }{ \cos (x) }\] do I get \[\cos ^{2} (x) + \sin ^{2} (x)\] Which is the Pythagorean identity and = 1?

OpenStudy (anonymous):

if you multiply all at the top you get cos(x)sin(x) and the bottom will give you sin(x)cos(x).......can also be written as cos(x)sin(x) so top and bottom will be the same

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No problem